# How quantum and evolutionary algorithms can help each other: two examples

### v1

@@ Acknowledgements
S.B. conducted the experiments and performed the simulations. All authors have analyzed the results and written the manuscript. S.N., S.D., and P.G.L. have revised the text and P.G.L. coordinated the overall project.
@@
The Generation class is another important class in the QUEVO module, responsible for managing the population of chromosomes. It is a collection of chromosomes that undergo evolution for a specified number of generations. After each evolution step, changes in the chromosomes in the generation occur by selecting a fixed number of chromosomes as elite and allowing the rest of them to evolve further. In each evolution step, the chromosomes are evaluated with the fitness function, and the fittest chromosomes become the parents for the next generation. The rest of the chromosomes are reset for the initial chromosomes. The Generation class stores a generation of chromosomes, the fitness associated with each chromosome, methods for running and retrieving fitness for two different fitness functions, and functions for printing. It provides methods for performing selection, crossover, and mutation to generate a new population of chromosomes.
@@
The compute_MW_entanglement method implements the Mayer-Wallach measure of entanglement, as a fitness function of the EA. This method processes a state vector representing a quantum circuit’s state, converting it into a $2^{n}$ x $2^{n}$ tensor, where $n$ is the circuit’s qubit count. The entanglement for each qubit $k$ is determined by tracing out all other qubits, summing the squared eigenvalues of their respective reduced density matrices. The results are then normalized by $1-\frac{1}{n}$, ensuring values fall between 0 and 1. More information and method usage are detailed in the QUEVO1 GitHub repository, accessible at https://github.com/shailendrabhandari/QUEVO1. Besides the well-known Mayer-Wallach measure, we evaluate the entanglement properties of quantum states using the von Neumann entropy as fitness functions. EA creates random density matrices from complex state vectors and calculates the von Neumann entropy based on their eigenvalues. This dual approach in measuring entanglement with Mayer-Wallach and von Neumann entropy enhances the quantum circuits’ entanglement characteristics, optimized via quantum evolutionary algorithms.
Supplementary Note II: State vector probabilities and density matrices for multi-qubit circuits
$P=|c|^{2}=a^{2}+b^{2},$ (11)
where $a$ represents the real part and $b$ the imaginary part of the coefficient. Following the computation of reduced density matrices for all qubits, the entanglement of the circuit is assessed using the Meyer-Wallach entanglement measure.
This appendix explicitly presents maximally entangled states for three, four, and five qubits. Results are systematically organized in tables, each corresponding to distinct quantum circuit configurations. Table 2 illustrates the outcomes for a three-qubit system, with reduced density matrices for circuits comprising 3 and 12 gates (Left), and their associated quantum state probabilities (Right). Table 3 explores a four-qubit system, detailing both reduced density matrices and state probabilities for four-gate circuits. Finally, Table 4 showcases results for a five-qubit system, encompassing reduced density matrices for circuits with five and twelve gates (Top), alongside the state probabilities (Bottom). These tables collectively provide a comprehensive view of quantum state entanglement across varying circuit complexities and gate counts.
@@
State Prob.3 Gates Prob.12 Gates
@@
Table 2: Final results for maximally entangled states with three qubits: (Left) Reduced density matrices for 3-gate and 12-gate three-qubit quantum circuits; (Right) Probabilities of quantum states for three-qubit circuits with three and twelve gates.
@@
$1$0.26709 0.16429 0.09833
Table 3: Final outcomes demonstrating maximally entangled states in a four-qubit system: (Left) Reduced density matrices for four-gate quantum circuits; (Right) Probabilities of quantum states for four qubit circuits with four gates.
@@
The density operator has several important properties. It is Hermitian, which means that it is equal to its conjugate transpose. It has a trace equal to one, reflecting the fact that it represents a pure state. Furthermore, it also satisfies the property, $\rho^{2}=\rho$, indicating the state is pure. In other words, the density matrix characterizes the pure state. To illustrate this, let us consider an example of a two-qubit, maximally-entangled pure state |$\ket{\psi_{AB}}$, which can be expressed as:
|$\ket{\psi_{AB}}=\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)=\frac{1}{\sqrt{2}}\begin{bmatrix}1\\ 0\\ 0\\ 1\end{bmatrix}.$ (14)
As an example, let’s reconsider the pure entangled state:
$\rho_{B}$ $=\text{Tr}_{A}(\rho_{AB})$
$=\frac{1}{2}\left[\text{Tr}_{A}(|0_{A}0_{B}\rangle\langle 0_{A}0_{B}|)+\text{Tr}_{A}(|0_{A}0_{B}\rangle\langle 1_{A}1_{B}|)+\text{Tr}_{A}(|1_{A}1_{B}\rangle\langle 0_{A}0_{B}|)+\text{Tr}_{A}(|1_{A}1_{B}\rangle\langle 1_{A}1_{B}|)\right]$
$=\frac{1}{2}\left[\text{Tr}(|0_{A}\rangle\langle 0_{A}|)|0_{B}\rangle\langle 0_{B}|+\text{Tr}(|0_{A}\rangle\langle 1_{A}|)|0_{B}\rangle\langle 1_{B}|+\text{Tr}(|1_{A}\rangle\langle 0_{A}|)|1_{B}\rangle\langle 0_{B}|+\text{Tr}(|1_{A}\rangle\langle 1_{A}|)|1_{B}\rangle\langle 1_{B}|\right]$
$=\frac{1}{2}\left[\langle 0_{A}|0_{A}\rangle|0_{B}\rangle\langle 0_{B}|+\langle 1_{A}|0_{A}\rangle|0_{B}\rangle\langle 1_{B}|+\langle 0_{A}|1_{A}\rangle|1_{B}\rangle\langle 0_{B}|+\langle 1_{A}|1_{A}\rangle|1_{B}\rangle\langle 1_{B}|\right]$
Starting with an entangled pure state $|\psi_{AB}\rangle$, the reduced density matrix $\rho_{B}$ of subsystem $B$ emerges as a mixed state. This matrix $\rho_{B}$ statistically represents subsystem $B$ by averaging out measurements from subsystem $A$. Tracing out $A$ yields $\rho_{B}$, reflecting the probability distribution of $B$. This approach is crucial for analyzing entangled states in complex systems, allowing insight into individual subsystems within entanglements. Key concepts and details about density matrices are sourced from [46, 42], especially the section "The density matrix and mixed states" (https://qiskit.org/textbook/ch-quantum-hardware/density-matrix.html##4.-The-Reduced-Density-Matrix--)
Supplementary Note IV: Evaluation of quantum purity and the Meyer-Wallach entanglement measure
In this appendix, a comprehensive derivation and analysis of the quantum purity of a state is provided, particularly focusing on its relation to the Meyer-Wallach measure. If the state |$\ket{\psi}$ is pure, then its density matrix is given by $\rho=\ket{\psi}\bra{\psi}$ |, and its purity yields:
[…]
The given equation is an expression for the quantum purity of a state |$\ket{\psi}$, denoted by $Q(\ket{\psi})$. The purity is a measure of how pure or mixed a quantum state is, and it is defined as the trace of the square of the density matrix $\rho$ that represents the state. For a pure state, the expression $1-\frac{1}{n}\sum_{k=0}^{n-1}Tr[\rho_{k}^{2}]$ is equal to zero. For a mixed state, the density matrix can be written as a convex combination of pure states, $\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ |, where $p_{i}$ are probabilities and $\sum_{i}p_{i}=1$. The purity of such a mixed state can not exceed $1$, as one can conclude from:
$\operatorname{Tr}[\rho^{2}]$ $=\operatorname{Tr}\left[\left(\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}\right)\left(\sum_{j}p_{j}\ket{\psi_{j}}\bra{\psi_{j}}\right)\right]$
[…]
$=\sum_{i}p_{i}^{2}+\sum_{i\neq j}p_{i}p_{j}|\braket{\psi_{i}}{\psi_{j}}|^{2}$

### v2

@@ Acknowledgements
S.B. conducted the experiments and performed the simulations. All authors have analyzed the results and written the manuscript. S.N., S.D., and P.G.L. have revised the text, and P.G.L. coordinated the overall project.
@@
The Generation class is another important class in the QUEVO module, responsible for managing the population of chromosomes. It is a collection of chromosomes that undergo evolution for a specified number of generations. After each evolution step, changes in the chromosomes in the generation occur by selecting a fixed number of chromosomes as elite and allowing the rest of them to evolve further. In each evolution step, the chromosomes are evaluated with the fitness function, and the fittest chromosomes become the parents for the next generation. The rate of the chromosomes is reset for the initial chromosomes. The Generation class stores a generation of chromosomes, the fitness associated with each chromosome, methods for running and retrieving fitness for two different fitness functions, and functions for printing. It provides methods for performing selection, crossover, and mutation to generate a new population of chromosomes.
@@
State Probability (3 Gates) Probability (12 Gates)
@@
Table 2: Final results for maximally entangled states with three qubits: (Top) Reduced density matrices for 3-gate and 12-gate three-qubit quantum circuits; (Bottom) Probabilities of quantum states for three-qubit circuits with three and twelve gates.
The compute_MW_entanglement method implements the MW measure of entanglement, as a fitness function of the EA. This method processes a state vector representing a quantum circuit’s state, converting it into a $2^{n}$ x $2^{n}$ tensor, where $n$ is the circuit’s qubit count. The entanglement for each qubit $k$ is determined by tracing out all other qubits, summing the squared eigenvalues of their respective reduced density matrices. The results are then normalized by $1-\frac{1}{n}$, ensuring values fall between 0 and 1. More information and method usage are detailed in the QUEVO1 GitHub repository, accessible at https://github.com/shailendrabhandari/QUEVO1. Besides the well-known MW measure, we evaluate the entanglement properties of quantum states using the von Neumann entropy as a fitness function. EA creates random density matrices from complex state vectors and calculates the von Neumann entropy based on their eigenvalues. This dual approach in measuring entanglement with MW and von Neumann entropy enhances the quantum circuits’ entanglement characteristics, optimized via quantum EAs.
Supplementary Note II: Evaluation of quantum purity and the MW entanglement measure
[…]
$=\sum_{i}p_{i}^{2}+\sum_{i\neq j}p_{i}p_{j}|\braket{\psi_{i}|\psi_{j}}|^{2}$
@@
$11$ -0.26709 0.16429 0.09833
Table 3: Final outcomes demonstrating maximally entangled states in a four-qubit system: (Top) Reduced density matrices for four-gate quantum circuits; (Bottom) Probabilities of quantum states for four-qubit circuits with four gates.
@@
[…]
As an example, let’s reconsider the pure entangled state:
$\rho_{B}$ $=\text{Tr}_{A}(\rho_{AB})$
$=\frac{1}{2}\left[\text{Tr}_{A}(|0_{A}0_{B}\rangle\langle 0_{A}0_{B}|)+\text{Tr}_{A}(|0_{A}0_{B}\rangle\langle 1_{A}1_{B}|)\right$.
$\quad+\left.\text{Tr}_{A}(|1_{A}1_{B}\rangle\langle 0_{A}0_{B}|)+\text{Tr}_{A}(|1_{A}1_{B}\rangle\langle 1_{A}1_{B}|)\right]$
$=\frac{1}{2}\left[\text{Tr}(|0_{A}\rangle\langle 0_{A}|)|0_{B}\rangle\langle 0_{B}|+\text{Tr}(|0_{A}\rangle\langle 1_{A}|)|0_{B}\rangle\langle 1_{B}|\right$.
$\quad+\left.\text{Tr}(|1_{A}\rangle\langle 0_{A}|)|1_{B}\rangle\langle 0_{B}|+\text{Tr}(|1_{A}\rangle\langle 1_{A}|)|1_{B}\rangle\langle 1_{B}|\right]$
$=\frac{1}{2}\left[\langle 0_{A}|0_{A}\rangle|0_{B}\rangle\langle 0_{B}|+\langle 1_{A}|0_{A}\rangle|0_{B}\rangle\langle 1_{B}|\right$.
$\quad+\left.\langle 0_{A}|1_{A}\rangle|1_{B}\rangle\langle 0_{B}|+\langle 1_{A}|1_{A}\rangle|1_{B}\rangle\langle 1_{B}|\right]$
Starting with an entangled pure state $|\psi_{AB}\rangle$, the reduced density matrix $\rho_{B}$ of subsystem $B$ emerges as a mixed state. This matrix $\rho_{B}$ statistically represents subsystem $B$ by averaging out measurements from subsystem $A$. Tracing out $A$ yields $\rho_{B}$, reflecting the probability distribution of $B$. This approach is crucial for analyzing entangled states in complex systems, allowing insight into individual subsystems within entanglements. Key concepts and details about density matrices are sourced from [10, 40], especially the section "The density matrix and mixed states" (https://qiskit.org/textbook/ch-quantum-hardware/density-matrix.html##4.-The-Reduced-Density-Matrix–)
Supplementary Note IV: State vector probabilities and density matrices for multi-qubit circuits
$P=|c|^{2}=a^{2}+b^{2},$ (20)
where $a$ represents the real part and $b$ the imaginary part of the coefficient. Following the computation of reduced density matrices for all qubits, the entanglement of the circuit is assessed using the MW entanglement measure.
This appendix explicitly presents maximally entangled states for three, four, and five qubits. Results are systematically organized in tables, each corresponding to distinct quantum circuit configurations. Table 2 illustrates the outcomes for a three-qubit system, with reduced density matrices for circuits comprising 3 and 12 gates (Top), and their associated quantum state probabilities (Bottom). Table 3 explores a four-qubit system, detailing both reduced density matrices and state probabilities for four-gate circuits. Finally, Table 4 showcases results for a five-qubit system, encompassing reduced density matrices for circuits with five and twelve gates (Top), alongside the state probabilities (Bottom). These tables collectively provide a comprehensive view of quantum state entanglement across varying circuit complexities and gate counts.